# Aerial Manipulation with Contact-Aware Onboard Perception and Hybrid Control

### v1

@@ IV-B Contact Factor
Concretely, when we switch to force holding mode, the robot is constrained along the wall normal direction. Between two consecutive states, the signed distance to the wall is preserved, and the velocity projected onto the wall normal is zero. These constraints are expressed as
@@ VII Conclusion
This paper developed a complete perception-control framework for aerial manipulation that eliminates reliance on an external motion capture system and enables robust contact-rich interactions. By incorporating contact-consistency factors into VIO, the estimator achieves improved accuracy during physical interaction, while the proposed IBVS and hybrid force-motion controller effectively regulate both contact wrenches and lateral motion. Experimental validation demonstrated accurate target approach, a 66.01% reduction in velocity estimation drift at contact, and stable force holding, all achieved using only onboard sensing. These results highlight the potential of our pipeline for more complex outdoor aerial manipulation tasks. Future work includes executing diverse aerial manipulation tasks under wind disturbance and in feature-sparse environments.

### v2

@@ IV-B Contact Factor
Concretely, the contact event is triggered by a force threshold. After we switch to force holding mode, the robot is constrained along the wall normal direction. Between two consecutive states, the signed distance to the wall is preserved, and the velocity projected onto the wall normal is zero. These constraints are expressed as
@@ VII Conclusion
This paper developed a fully onboard perception–control pipeline for contact-rich aerial manipulation without external motion capture system. By adding contact-consistency factors to VIO and integrating IBVS with a hybrid force–motion controller, the system improves estimation during physical interaction and regulates both contact wrench and lateral motion. Experiments demonstrate accurate target approach, stable force holding, and a 66.01% reduction in velocity estimation drift at contact with only onboard sensing, highlighting the potential for complex outdoor tasks. Future work will address online surface normal estimation and robustness to wind disturbance and feature-sparse environments.